OpenStudy (anonymous):

At the school carnival, Carmen sold 3 times as many hot dogs as Shawn. The two of them sold 152 hot dogs altogether. How many hot dogs did Carmen sell?

OpenStudy (anonymous):

You can use equation to solve this question.

OpenStudy (anonymous):

So suppose there Shawn sold x hotdogs. Then 3x+x=152 4x=152 x=152/4 x=38 3*38=114, so Carmen sold 114 hotdogs.

OpenStudy (texaschic101):

C + S = 152 C = 3S now sub 3S in for C in the first equation and solve for S (Shawns hotdogs) Once you find S, sub that back into either of the equations to find C (Carmen's hotdogs)